OpenStudy (anonymous):

binomial theorem (2x+4)^3

OpenStudy (ybarrap):

(2x+4)^3=(2x)^3 * 4^0+(2x)^2 * 4^1 + (2x)^1 * 4^2+ (2x)^0 * 4^3

OpenStudy (anonymous):

how is this exactly?

OpenStudy (ybarrap):

This is using the binomial theorem (left out the factorial terms): http://en.wikipedia.org/wiki/Binomial_theorem (2x+4)^3=(3C0)*(2x)^3 * 4^0+(3C1)*(2x)^2 * 4^1 + (3C2)*(2x)^1 * 4^2+ (3C3)*(2x)^0 * 4^3=8 x^3+48 x^2+96 x+64